### v1

# Mitigate the Gap: Investigating Approaches for Improving Cross-Modal Alignment in CLIP

Abstract
Contrastive Language–Image Pre-training (CLIP) has manifested remarkable improvements in zero-shot classification and cross-modal vision-language tasks. Yet, from a geometrical point of view, the CLIP embedding space has been found to have a pronounced modality gap. This gap renders the embedding space overly sparse and disconnected, with different modalities being densely distributed in distinct subregions of the hypersphere. In this work, we aim at answering two main questions: 1. Does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be mitigated by pushing apart the uni-modal embeddings via intra-modality separation? We design AlignCLIP, in order to answer these questions and show that answers to both questions are positive. Through extensive experiments, we show that AlignCLIP achieves noticeable enhancements in the cross-modal alignment of the embeddings, and thereby, reduces the modality gap, while maintaining the performance across several downstream evaluations, such as zero-shot image classification, zero-shot multi-modal retrieval and zero-shot semantic text similarity. The source code for reproducing our experiments is available at https://github.com/sarahESL/AlignCLIP.
1 Introduction
One of the most prominent and widely used pre-trained vision–language models is OpenAI’s Contrastive Language–Image Pre-training (CLIP) model (Radford et al., 2021). CLIP is a dual-stream vision–language encoder trained for learning a shared representation space, in which image and text modalities can be jointly embedded. It has demonstrated exceptional zero-shot capabilities for image classification, multi-modal retrieval as well as robustness to natural distribution shifts.
Figure 1: Modality gap and cross-modal alignment. (A) The average pairwise cosine similarity of the encoded image–text pairs is about 0.22, i.e., the average angle is about 78 degrees. (B) Schematic illustration of unaligned (left) versus aligned (right) embedding spaces.
Despite the outstanding performance of CLIP, recent work has shed light on a pronounced modality gap in the CLIP embedding space (Liang et al., 2022; Tyshchuk et al., 2023; Schrodi et al., 2024), leading to large distances between image and text embeddings. We illustrate these phenomena more comprehensively in Figure 1A, in which the DOSNES (Lu et al., 2019) projection of the CLIP-encoded image–text pairs from CC3M (Sharma et al., 2018) using the pre-trained ViT-B-32 backend is shown. As can be seen, each modality densely populates a separate small subregion of the CLIP’s representation embedding space. On the contrary, with a meaningful and well-structured cross-modal embedding space, we envision having similar samples fairly closely aligned with each other, regardless of their modalities. An example of such an embedding space is shown in Figure 1B, on the right side, with similar images and texts closely located on the hypersphere. In contrast, the left side of Figure 1B shows an example of an unaligned embedding space, inspired by the CLIP visualization.
Through extensive experiments, Liang et al. (2022) and Tyshchuk et al. (2023) showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Recently, Schrodi et al. (2024) showed that the driving factor behind the modality gap is the information-imbalance between the two modalities, i.e., texts in the training datasets usually have less detailed information in comparison to their pairing images. Furthermore, Jiang et al. (2023) show that exact modality alignment in multi-modal embeddings is suboptimal for downstream tasks.
Attempting to improve cross-modal alignment and reducing the modality gap in CLIP without losing performance in downstream tasks, we propose AlignCLIP, in the quest for answering two main questions: 1. Does sharing the parameter space between the multi-modal encoders reduce the modality gap? 2. Can the gap be further mitigated by pushing apart the uni-modal embeddings via intra-modality separation?
We answer the first question by sharing the transformer encoder and the projection layer in the vision and language encoders, and observe that it already results in noticeable cross-modal alignment improvements. As for the second question, we introduce the Intra-Modality Separation objective that encourages the embeddings within each modality to be expanded and pushed towards the opposite modality, i.e., text embeddings get pushed apart, towards image embeddings, and at the same, image embeddings get pushed apart and moved towards text embeddings. As a result, each modality gets moved towards the embeddings from the opposite modality and thereby, the modality gap can get reduced. Our experimental results support that the two aforementioned refinements show substantial improvement of the cross-modal alignment while maintaining, and slightly improving, the zero-shot transfer performance across a wide variety of downstream tasks.
2 Background, Notations and Concepts
Given a set of $N$ image–text pairs, we consider the CLIP image encoder to obtain the $l_{2}$-normalized vector $\vec{e}_{\mathrm{v}}^{\,i}\in\mathbb{R}^{d}$ as a $d$-dimensional embedding vector for image $v_{i}$ and the CLIP text encoder to obtain the $l_{2}$-normalized text embedding $\vec{e}_{\mathrm{t}}^{\,i}\in\mathbb{R}^{d}$ for the text sample $t_{i}$. We denote a batch of encoded image–text pairs by $E_{\mathrm{v}}\in\mathbb{R}^{b\times d}$ for images and $E_{\mathrm{t}}\in\mathbb{R}^{b\times d}$ for texts, where $b$ refers to the batch size.
Background on CLIP. The contrastive objective in CLIP is the average of the vision to language and the language to vision Info-NCE contrastive loss functions formulated as:
$\mathcal{L}_{v\rightarrow l}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{l\rightarrow v}=\frac{1}{N}\sum_{j=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,j}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}{\sum_{i=1}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}$ (1)
, respectively, where $\tau$ is the learnable temperature parameter. The overall CLIP loss is then:
$\mathcal{L}_{\textrm{clip}}=\frac{1}{2}\left[\mathcal{L}_{v\rightarrow l}+\mathcal{L}_{l\rightarrow v}\right].$ (2)
In practice, a symmetric cross-entropy loss is employed using the vision and language logits. The label $y\in\mathbb{R}$, which is the index of the image-text pair in the batch, represents the correspondence of the paired samples. For a batch of image–text pairs, $Y\in\mathbb{R}^{b}$ denotes the set of labels. Denoting $\langle A,B\rangle=A\cdot B^{\intercal}$, the visual and textual logits, $\hat{y}_{\mathrm{v}}\in\mathbb{R}^{b\times b}$ and $\hat{y}_{\mathrm{t}}\in\mathbb{R}^{b\times b}$, are then calculated as:
$\hat{y}_{\mathrm{v}}=\exp(\tau)\cdot\langle E_{\mathrm{v}},E_{\mathrm{t}}\rangle\quad,\quad\hat{y}_{\mathrm{t}}=\hat{y}_{\mathrm{v}}^{\intercal},$ (3)
respectively. Ultimately, the overall CLIP loss is calculated using the cross-entropy loss, $H$, as:
$\mathcal{L}_{\textrm{clip}}=\frac{1}{2}[H(\hat{y}_{\mathrm{v}},Y)+H(\hat{y}_{\mathrm{t}},Y)].$ (4)
Cross-Modal Alignment Score. In contrastive representation learning, the goal is to learn similar representations for positive pairs and distant representations for irrelevant negative samples. In cross-modal vision–language learning, often the paired image–texts form the positive pair distribution. Alignment entails mapping positive pairs to close embedding vectors such that a perfect alignment is achieved when $f(x_{1})=f(x_{2})$ for a given encoding function $f$ and a randomly drawn positive pair of samples $x_{1}$, $x_{2}$ (Wang and Isola, 2020). In this work, we are particularly interested in studying the alignment property for the CLIP embedding space as it represents the modality gap. We adopt the alignment measurement proposed by Goel et al. (2022) and define it as the average cosine similarity between the positive pairs in the CLIP embedding space, i.e., the paired image and text embeddings:
$\textrm{alignment}=\frac{1}{N}\sum_{i=1}^{N}\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i},\quad\quad\textrm{alignment}\in[-1,1].$ (5)
Higher scores demonstrate better alignment, and, therefore, a decrease in the modality gap.
3 AlignCLIP
Figure 2: Overview of sharing the transformer and projection layer in SharedCLIP.
3.1 Sharing the Learnable Parameter Space in CLIP
For answering whether a shared parameter space reduces the gap, AlignCLIP employs a transformer-based encoder architecture (Vaswani et al., 2017), where the transformer is shared between the two vision and language modality encoders. We suspect that one of the main reasons that the modality gap exists in the original CLIP’s embedding space is the fact that each modality has a separate disentangled set of parameters for optimization. As a result, even with the same architecture for encoding each of the modalities, the final values of the learned parameters associated to each modality can diverge quite radically. This difference can lead to distinct functions that map each modality to completely different subregions of the embedding space.
Therefore, we seek to align the outputs of the vision and language encoding functions by sharing their parameter space to the extent possible. Specifically, we share the parameters of the transformer encoder as well as the projection layer between the vision and language modalities. We utilize a standard transformer encoder architecture (Dosovitskiy et al., 2020). An overview of our refined overall model architecture is given in Figure 2. Yellow components are designated for the language modality, green components for the visual modality, and the blue parts are shared between the two modalities. For encoding texts, the yellow and blue parts of the model actively contribute in the embedding calculations. Similarly, when encoding images, the green components as well as the blue ones are invoked for the computation. Following the original CLIP, we use max-pooling for text embeddings and $[CLS]$ token embedding for the image embedding. For simplicity, we refer to this architecture as SharedCLIP throughout the rest of the paper. Sharing parameters in CLIP has been previously investigated from the perspective of downstream performance (You et al., 2022). In this work, we rather show its effectiveness with respect to the cross-modal alignment property.
3.2 Intra-Modality Separation
As shown in Figure 1A, each modality resides in a distinct dense subregion of the CLIP embedding space. We hypothesize that this phenomenon is a direct result of the cross-modal contrastive objective in CLIP, which merely optimizes the relative distances of image embeddings and text embeddings. Instead, a more practical scenario is to also consider the relative in-modality distances, i.e., pairwise distances of text embeddings and pairwise distances of image embeddings. We refer to this property as Intra-Modality Separation (IMSep) and define an objective function that enforces reasonable separation within each modality.
IMSep is achieved by a vision to vision as well as a text to text contrastive loss, where image–text pairs are the positive samples and any pairwise combination of image–image and text–text are considered as negative samples, respectively:
$\mathcal{L}_{v\rightarrow v}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1,\atop{i\neq j}}^{N}\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})\mathbin{/}\tau]},\quad\mathcal{L}_{t\rightarrow t}=\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp[(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,i})\mathbin{/}\tau]}{\sum_{j=1,\atop{i\neq j}}^{N}\exp[(\vec{e}_{\mathrm{t}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})\mathbin{/}\tau]}.$ (6)
In Figure 3, our approach for obtaining IMSep is schematically summarized. Given a batch of encoded image–text pairs, IMSep creates $\hat{y}_{\mathrm{vsep}}$ and $\hat{y}_{\mathrm{tsep}}$ logits and then minimizes the cross-entropy loss over $Y$ and $\hat{y}_{\mathrm{vsep}}$ as well as $\hat{y}_{\mathrm{tsep}}$.
To this end, first the pairwise cosine similarities are calculated by:
$\mathcal{V}=\langle E_{\mathrm{v}}\cdot E_{\mathrm{v}}\rangle,\quad\mathcal{T}=\langle E_{\mathrm{t}}\cdot E_{\mathrm{t}}\rangle,$ (7)
where $\mathcal{V}\in\mathbb{R}^{b\times b}$ and $\mathcal{T}\in\mathbb{R}^{b\times b}$. One should notice that while enforcing intra-modality separation, some samples might indeed be semantically similar to each other and therefore, must not be separated immensely. To this end, we regularize each intra-modality separation with respect to the pairwise semantic similarity of the samples. In order to calculate the semantic similarity within the uni-modal samples, we invoke a pre-trained sentence encoder and encode each text as $\vec{e}_{\mathrm{s}}^{\,i}\in\mathbb{R}^{d}$. We denote the corresponding batch of semantically encoded texts as $E_{\mathrm{s}}\in\mathbb{R}^{b\times d}$, and proceed to calculate the pairwise semantic similarity $S\in\mathbb{R}^{b\times b}$ and the distance $D\in\mathbb{R}^{b\times b}$ of the texts:
$\mathcal{S}=\frac{\langle E_{\mathrm{s}}\cdot E_{\mathrm{s}}\rangle}{\left\lVert E_{\mathrm{s}}\right\rVert^{2}},\quad\quad\mathcal{D}=1-\mathcal{S},$ (8)
respectively. We rely on $\mathcal{D}$ in order to re-scale $\mathcal{V}$ and $\mathcal{T}$. The goal of this re-scaling is to enforce a smaller dot product of the encoded images and texts, respectively, if they are semantically similar, i.e., have a smaller distance in $D$. Conversely, the dot products in $\mathcal{V}$ and $\mathcal{T}$ obtain larger values when the text samples are semantically distant according to $D$. By this re-scaling, we seek to control the minimization of $(\vec{e}_{\mathrm{v}}^{\,i}\cdot\vec{e}_{\mathrm{v}}^{\,j})$ as well as $(\vec{e}_{\mathrm{t}}^{\,i}\cdot\vec{e}_{\mathrm{t}}^{\,j})$ when the samples are semantically similar. This is because the re-scaling mechanism enforces reducing these in-modality dot product values, and therefore, the dot products do not get strongly minimized in the cross-entropy loss, since the values are already small. Re-scaling is performed by:
$\mathcal{V_{D}}=\mathcal{V}\odot\mathcal{D},\quad\mathcal{T_{D}}=\mathcal{T}\odot\mathcal{D}.$ (9)
where $\odot$ denotes the element-wise product, in which $\mathcal{V_{D}}\in\mathbb{R}^{b\times b}$ and $\mathcal{T_{D}}\in\mathbb{R}^{b\times b}$. Afterwards, we calculate $\mathcal{M}=\langle E_{\mathrm{v}},E_{\mathrm{t}}\rangle$ and mask the non-diagonal values by $\textrm{diag}(\mathcal{M})=\mathbb{I}\odot\mathcal{M}$. We then obtain the vision-vision logits $\hat{y}_{\mathrm{vsep}}\in\mathbb{R}^{b\times b}$ as well as text-text logits $\hat{y}_{\mathrm{tsep}}\in\mathbb{R}^{b\times b}$ by:
$\vspace*{-1mm}\hat{y}_{\mathrm{vsep}}=\exp(\tau)\cdot[\text{diag}(\mathcal{M})+\mathcal{V_{D}}],\quad\hat{y}_{\mathrm{tsep}}=\exp(\tau)\cdot[\text{diag}(\mathcal{M})+\mathcal{T_{D}}]$ (10)
Ultimately, we define the Intra-Modality Separation loss as:
$\mathcal{L}_{\textrm{IMsep}}=H(\hat{y}_{\mathrm{vsep}},Y)+H(\hat{y}_{\mathrm{tsep}},Y),$ (11)
and adopt the core of the CLIP loss to represent the cross-modal separation:
$\mathcal{L}_{\textrm{CRsep}}=H(\hat{y}_{\mathrm{v}},Y)+H(\hat{y}_{\mathrm{t}},Y).$ (12)
The final loss function optimized in AlignCLIP is:
$\mathcal{L}=\alpha\mathcal{L}_{\textrm{CRsep}}+\beta\mathcal{L}_{\textrm{IMsep}}$ (13)
Figure 3: Schematic summary of the Intra-Modality Separation approach in AlignCLIP.
4 Experiments
4.1 Training Dataset and Setup
We used the Conceptual Caption 12M (CC12M) dataset, which has been similarly used in previous work (Goel et al., 2022; Mu et al., 2022; Li et al., 2021; Changpinyo et al., 2021), for pre-training the models. We set our setup quite similar to that of the original CLIP with the ViT-B-16 backend, in order to ensure a fair comparison of CLIP with SharedCLIP and AlignCLIP. The pre-trained semantic encoder utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the SBERT all-mpnet-base-v2 model. In order to fairly compare the effectiveness of each model, we trained all of them from scratch using the CC12M dataset and the OpenCLIP implementation (Cherti et al., 2023; Ilharco et al., 2021). Each model was trained for 6 days using an NVIDIA H100 GPU with batch size 512 for 30 epochs using AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, a weight decay of $0.1$, and an initial temperature value of 0.07. The output embedding dimensionality for both the vision and language modalities is set to 768. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$. In terms of ablation study, we also trained IMSep loss without sharing the parameters between vision and language encoders. In Section A.1, more details about the training setup is provided. Furthermore, in Section A.3, we repeat our experiment when using output embedding dimension of 512 and show the results.
Model CC3M MSCOCO ImageNet-1K CIFAR-100 CIFAR-10
CLIP 0.42 0.47 0.41 0.38 0.4
SharedCLIP 0.59 0.62 0.57 0.54 0.54
AlignCLIP (no sharing) 0.67 0.71 0.66 0.59 0.61
AlignCLIP 0.63 0.67 0.62 0.58 0.6
Table 1: Comparison of the alignment score.
4.2 Cross-Modal Alignment
We start by reporting and comparing the alignment scores when using CLIP, SharedCLIP, and AlignCLIP models on the validation sets from CC3M, MSCOCO as well as the ImageNet-1K, CIFAR-100, and CIFAR-10 datasets. Table 1 summarizes the corresponding alignment scores. For ablations, we also report the results when the IMSep loss in AlignCLIP is trained without sharing the parameters. We observe that the original CLIP model has relatively low alignment scores, varying within $[0.38,0.47]$, across all five datasets. These scores mean that the average angle between the paired image–text embeddings is a value between 61 and 68 degrees. In contrast, sharing the parameter space in SharedCLIP results in noticeable improvements of up to 0.17 in the alignment scores. As a result, the average angle between the paired image–text embeddings decreases to about 51 degrees. Furthermore, using AlignCLIP yields even better alignment scores, ranging from 0.58 to 0.67, and a decreased average angle of 47 degrees between the multi-modal paired samples. Furthermore, AlignCLIP without sharing parameters achieves the best alignment scores. These observations confirm that AlignCLIP improves the cross-modal alignment in CLIP and thereby, reduces the modality gap.
We also plot the distribution of cosine similarities of the positive samples from MSCOCO validation dataset when encoded using the original CLIP, SharedCLIP, and AlignCLIP in Figure 4. We find that using SharedCLIP noticeably shifts the distribution of the cosine similarity of positive samples towards higher similarity values. A higher similarity of positive samples, i.e., image–text pairs, means achieving greater cross-modal alignment, and thus a lower modality gap. Furthermore, AlignCLIP shifts the distribution even more to the right, resulting in higher similarities of positive samples, better cross-modal alignment, and a reduction of the modality gap.
Figure 4: Pairwise cosine similarities of positive samples in MSCOCO.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 31.4 58.7 28.1 55.9 61.5 95.6 35.2 63.3 11.6 36.5
SharedCLIP 32.1 59.7 26.4 54.7 56.9 95.2 32.8 61.2 10.7 35.4
AlignCLIP (no sharing) 28.7 55.8 23.1 49.7 50.1 91.4 33.1 61.8 8.4 30.7
AlignCLIP 32.4 60 31.2 61.8 66.4 96.9 33.1 63.3 11.8 37.1
Table 2: Accuracy scores for zero-shot image classification.
4.3 Zero-Shot Transfer
CLIP’s pre-training objective for predicting whether a text is paired with an image has resulted in outstanding zero-shot image classification and multi-modal retrieval results (Radford et al., 2021). Therefore, we further assess how sharing the learnable parameters and the intra-modality loss affect the zero-shot transfer to donstream tasks. Furthermore, Section A.2 reports the results for image classification with linear probing.
Image Classification. We conduct the zero-shot classification experiments on ImageNet-1K (Russakovsky et al., 2015), CIFAR-100, CIFAR-10 (Krizhevsky et al., 2009), Food-101 (Bossard et al., 2014), and Stanford Cars (Krause et al., 2013) with the combination of text prompts used by CLIP (Radford et al., 2021), e.g., “a photo of the {label}”. The experimental results summarized in Table 2 show that SharedCLIP reduces the accuracy of the zero-shot classification on CIFAR-10 by about $5\%$ and $0.4\%$ when measuring Top-1 and Top-5 accuracy scores, respectively. Similarly, SharedCLIP’s results on CIFAR-100 show about $2\%$ and $1\%$ decrease in accuracy. The trend of accuracy reduction is also observed on Food-101 and Stanford Cars datasets. However, on ImageNet-1K, SharedCLIP evinces up to $1\%$ improvement of accuracy. In contrast, AlignCLIP yields the best scores across all five datasets. On CIFAR-10, AlignCLIP achieves up to $5\%$ and $1\%$ improvements in Top-1 and Top-5 accuracy scores, when compared to the original CLIP model. Similarly, using AlignCLIP for CIFAR-100 results in about $2\%$ and $6\%$ enhancement in Top-1 and Top-5 accuracy scores in comparison to the original CLIP. The trend of improvement is also observed on the Stanford Cars dataset. Furthermore, AlignCLIP achieves $1\%$ and $1.3\%$ improvement of accuracy on ImageNet-1K when compared to the original CLIP. On the Food-101 dataset, the original CLIP model achieves the best Top-1 accuracy of 35.2 and AlignCLIP achieves comparable results in Top-5 accuracy. According to our ablation study, AlignCLIP without sharing the parameters between encoders results in a noticeable decrease of the classification performance. Our experiments thus evince that via sharing parameters and the additional intra-modality separation, AlignCLIP improves the cross-modal alignment of the embeddings without losing performance on the downstream zero-shot image classification task.
Multi-Modal Retrieval. In addition to classification, we evaluate SharedCLIP and AlignCLIP in the application of zero-shot image-to-text and text-to-image retrieval using the validation splits from the MSCOCO (Lin et al., 2014) and Flickr30K (Plummer et al., 2015) datasets. The results of these evaluations are reported in Table 3. In all settings, the text prompt “a photo of the {caption}” is used. Our experiments show that both SharedCLIP and AlignCLIP improve the retrieval results measured by $R@\{1,5,10\}$ on both datasets when compared to the original CLIP model. In addition, AlignCLIP achieves comparable results to SharedCLIP in $R@\{5,10\}$. In summary, the best results for image-to-text and text-to-image retrieval on both MSCOCO and Flickr30K datasets are achieved by SharedCLIP, with AlignCLIP ranking in a very close second position. Finally, according to our ablation, AlignCLIP without sharing the parameters between encoders decreases the retrieval performance in comparison to the original CLIP model. These experiments demonstrate that it is possible to reduce the modality gap in CLIP via parameter sharing while improving the downstream multi-modal retrieval tasks. Furthermore, the addition intra-modality separation improves the alignment noticeably without losing perofamnce when compared to the original CLIP model.
MSCOCO Flickr30K
Model I $\rightarrow$ T T $\rightarrow$ I I $\rightarrow$ T T $\rightarrow$ I
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 31.4 57 68.6 20.5 44.1 55.9 53.2 80.5 88.6 39.9 69 78.5
SharedCLIP 33.5 59.6 70.8 21.8 45.4 57.3 58.3 83.6 89.8 42.5 70 79.1
AlignCLIP (no sharing) 26.2 52 63.8 17.8 40.1 52.2 48.5 74.6 82.7 35.5 63.2 74
AlignCLIP 32.7 59.2 70.8 21.1 44.6 56.3 54.8 82.7 90.4 40.5 69.8 79.1
Table 3: Zero-shot cross-modal retrieval summarized with R@{1, 5, 10}.
Semantic Text Similarity. We further evaluate SharedCLIP and AlignCLIP models on seven Semantic Textual Similarity (STS) benchmarks using the SemEval toolkit (Conneau and Kiela, 2018). Measuring the level of semantic similarity between encoded texts using cosine similarity, we report the Pearson correlations between the human-assigned similarity scores, provided as the ground-truth in the STS benchmark datasets, and the cosine similarity of the encoded texts in Table 4. SharedCLIP shows slight improvements on STS14 and STS15 while reducing the correlation on the rest of the benchmarks. This shows that sharing the transformer between the image and text modalities damages the semantic representation of the texts. In contrast, the intra-modality separation with the semantic similarity re-scaling mechanism in AlignCLIP substantially corrects this damage and improves the results noticeably. As a result, AlignCLIP achieves the best average results on the STS task. Similar to the previous tasks, AlignCLIP without sharing parameters substantially decreases the performance. These observations consistently show that AlignCLIP, while reducing the modality gap, achieves comparable results to CLIP in the zero-shot evaluation of semantic text similarity.
Model
STS12
STS13
STS14
STS15
STS16
STS-B
SICK-R
Average
CLIP 0.55 0.42 0.54 0.59 0.60 0.59 0.79 0.58
SharedCLIP 0.55 0.40 0.55 0.60 0.58 0.59 0.72 0.57
AlignCLIP (no sharing) 0.52 0.36 0.53 0.58 0.60 0.58 0.78 0.56
AlignCLIP 0.56 0.43 0.60 0.64 0.64 0.59 0.77 0.6
Table 4: Pearson correlations for zero-shot STS using the text embeddings.
4.4 Robustness to Natural Distribution Shift
In the zero-shot image classification task, CLIP has additionally shown impressive robustness to natural distribution shifts and promising generalizability to out-of-distribution images. Therefore, we expand our evaluations and investigate to what extent SharedCLIP and AlignCLIP change the performance with natural distribution shifts. We use the ImageNetV2, ImageNet-R, ImageNet-A, ImageNet-O and ImageNetSketch datasets for these evaluations and report the corresponding results in Table 5, in terms of Top-1 and Top-5 accuracy. In summary, SharedCLIP achieves the best results on ImageNet-R, ImageNet-O and ImageNetSketch, whereas the best results on ImageNetV2 and ImageNet-A are achieved by AlignCLIP. Based on these observations, we conclude that sharing parameters and applying intra-modality separation in SharedCLIP and AlignCLIP preserve the robustness to natural distribution shifts when compared to the original CLIP model and at the same time, improve the cross-modal alignment. However, the IMSep objective in AlignCLIP without sharing parameters of the encoders results in a decrease in the classification accuracy.
Model ImageNetV2 ImageNet-R ImageNet-A ImageNet-O ImageNetSketch
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 27.1 53.3 39.8 65.8 6.5 25.4 25.8 45.5 19.4 41.8
SharedCLIP 27.5 53.5 40.2 67.3 6.7 25.5 26.2 45.8 20.6 43.2
AlignCLIP (no sharing) 24.6 50 34.5 59.7 5.8 23.2 23.2 42 17.6 38.6
AlignCLIP 27.7 53.9 38.6 65 6.9 25.5 24.8 45.8 19.7 42
Table 5: Accuracy scores for zero-shot classification and natural distribution shift.
4.5 DOSNES Visualizations
For a more comprehensive comparison of the distribution of each modality, we visualize the DOSNES projection of the encoded image–texts from the MSCOCO validation set in Figure 5. As can be seen, the uni-modal embeddings in CLIP are densely located on opposite sides of the hypersphere. In contrast, the embeddings get spread out when using the SharedCLIP model. Finally, the embedding space of AlignCLIP achieves the best spread of the uni-modal embeddings and substantially greater alignment of image and text embeddings. Thus, the intra-modality separation leads to a better alignment and a substantial reduction of the modality gap. Furthermore, Figure 5 shows that AlignCLIP reduces the sparsity of the embeddings on the hypersphere.
Figure 5: DOSNES visualization of the multi-modal embeddings using CC3M.
5 Related Work
Modifications of CLIP. Prior work has sought to improve several different aspects of CLIP. UniCLIP (Lee et al., 2022) proposed a multi-positive contrastive loss for training augmentation-aware feature embeddings. SLIP (Mu et al., 2022) added an additional image-based contrastive objective function using augmented images as positive samples to the original CLIP’s loss function. Furthermore, CyCLIP (Goel et al., 2022) formalized the geometric consistency of the image and text embeddings in the CLIP embedding space and added a loss function for regularizing the cross-modal and in-modal similarity scores. The authors provide an analysis of the cross-modal alignment showing that the final CyCLIP model does not improve the alignment of image–text embeddings. In xCLIP (Zhou et al., 2023), a non-contrastive training regimen based on image–text pairs is adopted in order to mitigate the requirement of large batch sizes in contrastive learning. With a similar goal, SigLIP (Zhai et al., 2023) employed a pairwise sigmoid loss for scaling up the batch size. The recently published ReCLIP (Hu et al., 2024) had the goal of refining the fine-tuning of CLIP in domain adaptation settings, based on unlabeled target samples. While ReCLIP results in noticeable gains for the considered classification task, its effectiveness with regard to the cross-modal alignment property is not studied by the authors. In contrast to previous work, AlignCLIP studies the cross-modal alignment and modality gap phenomena in CLIP by investigating the effects of parameter-sharing as well as intra-modality separation.
Modality Gap in CLIP. The modality gap in CLIP’s embedding space was first studied by Liang et al. (2022). Authors showed that the modality gap is caused by a combination of the model initialization and the contrastive loss optimization. Furthermore, they showed that CLIP’s embedding space is very sparse such that the effective embedding space is an extremely narrow cone. In a similar study, Tyshchuk et al. (2023) measured the alignment of image and text embeddings in CLIP with a focus on the isotropic properties. More recently, Schrodi et al. (2024) showed that one of the key factors contributing in both the modality gap and the bias in CLIP is the information-imbalance between the two modalities, i.e., images often have much more detailed information of the scene in comparison to their corresponding text descriptions. Furthermore, Jiang et al. (2023) showed that exact modality alignment in multi-modal embeddings is suboptimal for downstream tasks and propose intra-modality as well as inter-modality regularization using augmented samples in order to improve the latent embeddings structures. In contrast to the previous work, we study the effectiveness of parameter sharing between the modality encoders as well as semantically-regularized separation of the uni-modal embeddings on reducing the modality gap.
6 Conclusion
This work investigates the potential of reducing the modality gap in CLIP by sharing the learnable parameter space of the vision and language encoders as well as enforcing a contrastive intra-modality separation objective. We further examine the effects of modality gap reduction by the aforementioned refinements in the performance of downstream tasks. Through extensive experiments, we show that SharedCLIP and AlignCLIP improve various zero-shot downstream applications when compared to CLIP, while substantially improving the cross-modal alignment, and therefore, reducing the gap. Furthermore, our ablation on the intra-modality loss shows that sharing parameters between the vision and language encoders is necessary for preventing decrease in the performance of the downstream tasks. Our work shows that it is possible to mitigate the modality gap in CLIP via parameter sharing and intra-modality separation without losing downstream performance.
Acknowledgments
We acknowledge the financial support from the German Federal Ministry for Education and Research (BMBF) within the project KI-Servicezentrum Berlin Brandenburg (01IS22092).
Appendix A Appendix / Supplemental Material
A.1 Training Dataset and Setup
For pre-training, we used the Conceptual Caption 12M (CC12M) dataset [Changpinyo et al., 2021]. In comparison to the data used for pre-training CLIP, i.e., 400M image–text pairs, CC12M is a much smaller dataset, with about 12M noisy image–text pairs. Therefore, our results are not directly comparable to the numbers reported in the original CLIP paper [Radford et al., 2021]. Nonetheless, CC12M is one of the popularly used large-scale and publicly available datasets that enables pre-training vision–language models and analyzing the effectiveness of different training paradigms on benchmark evaluations.
We adopted a transformer encoder consisting of 12 layers and 12 heads in SharedCLIP and AlignCLIP. The same input pre-processing and augmentations employed in CLIP have been used for SharedCLIP as well as AlignCLIP, including random cropping of images to the size $224\times 224$. The image patch size for encoding visual data is set to $16\times 16$. When encoding texts, the maximum sequence length is set to 77 tokens and the vocabulary size for the embedding layer is set to 49,408. The output embedding dimensionality for both the vision and language modalities is set to 768. We chose our setup to be quite similar to that of the original CLIP with the ViT-L-16 backend, in order to ensure a fair comparison of CLIP and our proposed model. SBERT all-mpnet-base-v2 model has been utilized in AlignCLIP for re-scaling text–text and image–image cosine similarities is the .
In order to fairly compare the effectiveness of each model, we trained all of the models , i.e., the original CLIP with ViT-B-16 backend, SharedCLIP, and AlignCLIP, from scratch using the CC12M dataset and the OpenCLIP implementation [Cherti et al., 2023, Ilharco et al., 2021]. For all models, we used AdamW optimization with a starting learning rate of $1\times 10^{-3}$, cosine scheduler, 10,000 warmup steps, and a weight decay of $0.1$. The initial temperature value for all models were set to 0.07. Each model was trained using an NVIDIA H100 GPU with batch size 512 for 30 epochs. We used the checkpoint from the last epoch in our evaluations of downstream experiments. In AlignCLIP, we set $\alpha=1$ and $\beta=\frac{1}{2}$.
A.2 Linear Probing for Image Classification
Table 6 shows the results of image classification with linear probing.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
CLIP 50 62.6 85 67 42.2
SharedCLIP 51.2 63 85 68.3 40.5
AlignCLIP 50.3 63.4 85.8 67.9 42.2
Table 6: Accuracy scores for image classification with linear probing.
A.3 Experimental Results With Lower Output Dimensions
In order to examine how the trend of our observations change based on the choice of output embedding dimension, we repeated our training and downstream tasks using the output dimension of 512. We used the same hyperparatmer setting described in Section A.1. The corresponding results are reported in Table 7 to Table 9
Model CC3M MSCOCO ImageNet-1K CIFAR-100 CIFAR-10
CLIP 0.42 0.47 0.41 0.38 0.4
SharedCLIP 0.6 0.64 0.59 0.55 0.57
AlignCLIP 0.63 0.68 0.63 0.58 0.59
Table 7: Comparison of alignment scores when output dimension 512.
Model ImageNet-1K CIFAR-100 CIFAR-10 Food-101 Stanford Cars
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
CLIP 31.5 59 27.6 57.1 65.6 95.6 33.2 61.9 11.6 37.4
SharedCLIP 31.6 59 29 58.7 59.5 95.6 33.7 62.5 9.4 32.7
AlignCLIP 30.9 58.6 26.5 53.7 53.3 92.8 31.7 60.5 9.7 33.4
Table 8: Accuracy scores for zero-shot image classification when output dimension 512.
MSCOCO Flickr30K
Model I $\rightarrow$ T T $\rightarrow$ I I $\rightarrow$ T T $\rightarrow$ I
R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
CLIP 31.7 57.9 68.9 20.5 44.3 56 54.5 78.7 87.3 40.3 68.6 78.8
SharedCLIP 32.8 59.3 70.4 21 44.6 56.6 56.1 81.5 87.3 41.2 69.5 78.9
AlignCLIP 30.6 56.8 68.5 20.3 43.6 55.4 54.6 80.5 88 40.6 68.1 77.5
Table 9: Zero-shot cross-modal retrieval summarized with R@{1, 5, 10} when output dimension 512.
Appendix B Limitations
For the zero-shot classification experiments, we have tested the models using the ImageNet-1k, CIFAR-10, CIFAR-100, Stanford Cars, and Food-101 datasets. More classification analysis using other benchmarks datasets such as Flower-102 and Hatefule Memes has not been studied in this work. Furthermore, this study is limited to the English language and further analyses on multilingual and cross-lingual representation learning are necessary. Moreover, the re-scaling mechanism in the intra-modality separation loss is dependant on the choice of pre-trained sentence encoder and has not been fully benchmarked in this work.
Appendix C Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning at a fundamental level with regard to cross-modal representation learning. We acknowledge that this line of work has a broad range of potential societal implications. For instance, vision–language models may exhibit harmful biases and stereotypes, particularly when trained on data crawled from the Web. Due to their incorporation in prominent generative AI models, vision–language models may also contribute toward the model’s ability to produce images portraying trademarked characters or notable figures. These sorts of concerns need to be carefully considered before incorporating such models into real-world applications.